A manufacturer in a monopolistically competitive industry produces two different brands of a product for which the demand functions are P1=56-4Q1 and P2=48-2Q2. and the joint cost function is



TC=Q12+5Q1Q2+Q22. Find the profit maximizing level of output and the price that should be charged for each brand.

1
Expert's answer
2022-01-17T12:36:56-0500

"P_1= 56-4Q_1"

"P_2= 48-2Q_2"

"TC=Q_1^2+5Q_1Q_2+Q_2^2"

"\Pi=TR-TC"

"= P_1Q_1+P_2Q_2" - "(Q_1^2+5Q_1Q_2+Q_2^2)"

"=Q_1(56-4Q_1)+Q_2(48-2Q_2)-Q_1^2-5Q_1Q_2-Q_2^2"

="56Q_1-4Q_1^2+48Q_2-2Q_2^2-Q_1^2-5Q_1Q_2-Q_2^2"

First Order Condition

"\frac{\Delta\Pi}{\Delta Q_1}=56-10Q_1-5Q_2=0"

"\frac{\Delta\Pi}{\Delta Q_2}=48-6Q_2-5Q_1=0"

Solve the two Equations Simultaneously by eliminating "Q_1"

"56-10Q_1-5Q_2=0"

"(48-6Q_2-5Q_1=0)2"

="56-10Q_1-5Q_2=0"

"96-10Q_1-12Q_2=0"

40=7Q"_2"

"Q_2= 5.7"

Eliminate Q"_2 from the above equations"

="(56-10Q_1-5Q_2=0)6"

"=(48-6Q_2-5Q_1=0)5"

"336-60Q_1-30Q_2=0"

"240-30Q_2-25Q_1= 0"

"35Q_1=96"

"Q_1= 2.7"

From the Inverse Demand functions

"P_1=56-4(2.7)=45.2"

"P_2=48-2(5.7)=36.6"